OpenStudy (dmezzullo):

Create your own two-step equation with NO SOLUTION as the solution. Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS.

OpenStudy (dmezzullo):

@jazy

OpenStudy (dmezzullo):

@mikala1

OpenStudy (dmezzullo):

@jhonyy9

OpenStudy (anonymous):

ghold on ik the answer

OpenStudy (dmezzullo):

k

OpenStudy (anonymous):

7t+3=7t-3 7t-7t+3=7t-7t-3 3= -3 <- answer

OpenStudy (anonymous):

BECAUSE is not a correct statement, there is no solution for your original equation. In fact, it represents 2 parallel lines, so they never intersect at a commoin point.

OpenStudy (dmezzullo):

tyvm medal?

OpenStudy (dmezzullo):

bye guysh gtg bike